# Many-body enhancement of energy storage in a waveguide-QED quantum battery

### v2

@@ Abstract
Discussion.— We have analyzed two different arrangements of a one-dimensional waveguide-QED system: an ordered and a disordered array. We performed extensive numerical simulations of the time evolution for these two systems and compared the energy decay to a single-atom benchmark. In the ordered case, the protection properties were enhanced by many-body effects of the system, but only at specific lattice spacings where the optical lattice behaves as an insulator. In contrast, in the disordered case, we discovered that, regardless of the parameters chosen, the system protects a fraction of the energy and ergotropy within the lattice for extended periods. Therefore, we demonstrated how collective waveguide-QED effects can enhance energy storage performances. In the future, it would be tempting to experimentally test our predictions using atoms in a waveguide Roy et al. (2017); Chang et al. (2018) or with molecular excitons Tibben et al. (2024). A QB is a complex quantum device where energy can be injected, stored, and finally extracted. Therefore, it is important to interface our findings, where collective effects enhance energy storage, with cases where collective effects are used to boost charging Quach and Munro (2020); Rossini et al. (2020); Gyhm et al. (2022). Additionally, energy extraction should be considered in this context.
@@ Acknowledgements.
Acknowledgments.— We thank M. Polini and D. E. Chang for useful discussions and insightful suggestions. S.T. acknowledges financial support by MUR (Ministero dell’ Università e della Ricerca) through the following projects: PNRR MUR project PE0000023-NQSTI, PRIN 2017 Taming complexity via Quantum Strategies: a Hybrid Integrated Photonic approach (QUSHIP) Id. 2017SRN-BRK, and project PRO3 Quantum Pathfinder. G.M.A. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101002955 – CONQUER). G.C. acknowledges that results incorporated in this standard have received funding from the T-NiSQ consortium agreement financed by QUANTERA 2021 and that this work is supported in part by the Italian MUR Departments of Excellence grant 2023-2027 “Quantum Frontiers”.
SUPPLEMENTAL MATERIAL
Full Hamiltonian
The general Hamiltonian of the system is:
[…]
$+g\sum_{j,\nu}\int_{-\infty}^{\infty}\left(b_{\nu,k}\sigma_{eg}^{j}e^{i\nu kz_{j}}+{\rm h.c.}\right)\,dk,$
where $H_{0}^{L}=\omega_{0}\sum_{j=1}^{L}(I-\sigma_{z}^{j})$ is the sum of the local Hamiltonians acting on the two-level systems (TLSs), or atoms [$I$ is the $2\times 2$ identity matrix and $\sigma_{z}^{j}$ the $Z$ Pauli matrix acting on the $j$th site], with $\omega_{0}$ the energy of the excited state of the atoms. The $z_{j}$ variables denote the atom positions, $g$ is the coupling constant between the electromagnetic field and the artificial atoms of the optical lattice, which depends on the dipole moment of the photonic transition, $\sigma_{eg}^{j}$ is the operator which takes the excited state of the $j$th TLS and brings it to the ground state, finally $b_{\nu,k}^{(\dagger)}$ is the annihilation (creation) operator acting on the electromagnetic field with wavevector $k$ at frequency $\omega_{k}$, while $\nu={\pm}$ is the propagation direction of the electromagnetic mode. In this general Hamiltonian, we are considering a continuum of modes.
After employing a Born-Markov approximation, the photonic degree of freedom can be integrated out, such that the model is effectively described by the Lindblad master equation given in Eq. (1) of the main text Sheremet et al. (2023).
This model has been analyzed extensively in the literature Chang et al. (2012); Dzsotjan et al. (2010); Asenjo-Garcia et al. (2017); Lalumière et al. (2013); Albrecht et al. (2019).
Appendix A Single atom benchmark
We consider an initially excited atom, $\rho(0)=\ket{1}\!\!\bra{1}$ |, decaying into the one-dimensional electromagnetic environment at a rate that depends on the wavevector $k$, i.e., $\Gamma_{0}=\Gamma_{1D}[1-\cos(k_{1D}|2d|)]$. The time evolution of its reduced density operator reads
$\rho(t)=\begin{pmatrix}1-e^{-\Gamma_{0}t}&0\\ 0&e^{-\Gamma_{0}t}\end{pmatrix}.$ (8)
$\text{Tr}\big{[}\rho(t)\,H_{0}^{1}\big{]}=\omega_{0}e^{-\Gamma_{0}t},$ (9)
and the corresponding ergotropy, that reads:
$\mathcal{E}[\rho(t)]=\left\{\begin{array}[]{cc}\omega_{0}(2e^{-\Gamma_{0}t}-1)&\;\;{\rm if}\;\;t<\log 2/\Gamma_{0},\\ 0&\;\;{\rm if}\;\;t\geq\log 2/\Gamma_{0}.\end{array}\right.$ (10)
Figure 4: (a) The temporal decay rate $\gamma$ of the energy at long times, in the ordered scenario. Orange circles denote the numerical results of an exponential fit for the $E_{M,L}(\tau)$ data at fixed $L$, while the straight blue line represents the power-law fitting function $\gamma=A/L^{3}$, with $A\approx 0.747$. (b) The temporal decay rate $\gamma_{\mathcal{E}}$ of the ergotropy. Orange circles are obtained by exponentially fitting the $\mathcal{E}_{M,L}(\tau)$ data at fixed $L$, while the straight blue line is a power-law fit $\gamma_{\mathcal{E}}=B/L^{6}$, with $B\approx 1139.2$. In both panels we use a log-log scale, set $k_{1D}d=2.7\pi$, $M=3$, and the quantities $E_{M,L}(\tau)$ and $\mathcal{E}_{M,L}(\tau)$ are rescaled for the input energy of the QB. In all plots both $\gamma$ and $\gamma_{\mathcal{E}}$ are computed in units of $\Gamma_{1D}$.
Decay rates
In this section we focus on the behavior of the decay rates for the energy ($\gamma$) and the ergotropy ($\gamma_{\mathcal{E}}$) at long times, for the ordered model. As reported in the main text, in the ordered case, both the energy and the ergotropy eventually follow an exponential time behavior: the decay rate thus depends on the parameter $k_{1D}$, the lattice-spacing $d$, and the system size $L$. Concerning the decay of the energy $E_{M,L}(\tau)$, Fig. 4(a) shows that the dependence of $\gamma$ on the $L$ is $\gamma\propto L^{-3}$, in accordance with the existing literature Albrecht et al. (2019); Henriet et al. (2019); Zhang and Mølmer (2019).
Conversely, in Fig. 4(b) we present a novel result: the ergotropy in the ordered model seems to fall down much faster than the energy: in fact, by a fit of the numerical data, we extrapolate a behavior $\gamma_{\mathcal{E}}\propto L^{-6}$. Even though our numerics believably suggests this scaling, we are only able to explore relatively small system sizes, thus we cannot definitely rule out the robustness of this behavior in the thermodynamic limit.
Figure 5: (a) The local energy $E_{\{j\},L}(\tau)$, for each site $j$ of the chain, as a function of the evolution time $\tau$, in the ordered case. (b) The mean local energy $\langle\!\langle E_{\{j\},L}(\tau)\rangle\!\rangle$, after averaging over $10^{4}$ realizations, for each site $j$ of the chain, as a function of $\tau$, in the disordered case. In both panels we consider a system of size $L=9$ and load an initial state with $M=3$ perfectly charged atoms nearest to the mirror. In all plots the local input energy of each charged site is $E_{\{j\},L}(0)=1$ for $j=1,2,3$, and we fix $k_{1D}d=2.7\pi$.
Energy localization
In this section we show the temporal behavior of the average local energy for each TLS of the lattice, defined as
$E_{\{j\},L}(\tau)=\text{Tr}\big{[}\rho_{j}(\tau)\,(I-\sigma_{z}^{j})\big{]},$ (11)
where $\rho_{j}(\tau)={\rm Tr}_{L-\{j\}}[\rho(\tau)]$ is the reduced density matrix corresponding to the $j$th site. Figure 5(a) shows the local energy behavior in time of the ordered system; here we simulate a system of size $L=9$ and with the first $M=3$ sites initially charged. We see that the discharged sites pick up energy over time and the system eventually loses energy exponentially in time. In contrast, Fig 5(b) reports the results for the disordered system (we choose again $L=9$ and $M=3$). At long times, the first three sites retain a larger fraction of the input energy compared to the ordered case, while the other sites do not get fully charged in time.
These findings confirm that, in the disordered case, a fraction of the energy remains localized in the charged sites freezing the energy dynamics in time, thus they are in substantial agreement with the many-body localization results of Ref. Fayard et al. (2021).

### v3

@@ III Results
IV Discussion
We have analyzed two different arrangements of a one-dimensional waveguide-QED system: an ordered and a disordered array. We performed extensive numerical simulations of the time evolution for these two systems and compared the energy decay to a single-atom benchmark. In the ordered case, the protection properties were enhanced by many-body effects of the system, but only at specific lattice spacings where the optical lattice behaves as an insulator. In contrast, in the disordered case, we discovered that, regardless of the parameters chosen, the system protects a fraction of the energy and ergotropy within the lattice for extended periods. Therefore, we demonstrated how collective waveguide-QED effects can enhance energy storage performances. Let us stress that in this study we analyzed a waveguide instead of a cavity, because on the one hand it gives a channel to charge the battery without changing the structure of the system. On the other hand, it models the interaction between the battery and the environment more realistically; in fact, in this model, we consider energy leaking from the lattice boundary. We notice, however, that these many-body effects may be detrimental in the charging and discharging phase, because they induce a strong energetic rigidity in the system, which is essential during the storage, but it might be a hindrance to change the charge of the QB. Thus, a fully operational QB should ideally exhibit two distinct operational phases: one dedicated to the efficient charging and discharging, and another one optimized for the long-term energy storage.
In the future, it would be tempting to experimentally test our predictions using atoms in a waveguide Roy et al. (2017); Chang et al. (2018) or with molecular excitons Tibben et al. (2024). In realistic implementations, the coupling to modes outside the waveguide would rapidly compromise the QB storage capacity. While this issue is relatively less important in circuit-QED platforms Astafiev et al. (2010); Mirhosseini et al. (2019); Kannan et al. (2023); Brehm et al. (2021), it becomes critical in the optical regime, where only a small fraction of the atomic emission couples into the waveguide. One potential solution to this challenge involves using sub-wavelength atomic arrays, where collective free-space emission could be engineered to preferentially radiate into the waveguide via the so called “selective radiance” mechanism Asenjo-Garcia et al. (2017). A QB is a complex quantum device where energy can be injected, stored, and finally extracted. Therefore, it is important to interface our findings, where collective effects enhance energy storage, with cases where collective effects are used to boost charging Quach and Munro (2020); Rossini et al. (2020); Gyhm et al. (2022). In particular, we remark that waveguide-QED systems are capable of being used in any of the phases of a QB. In fact, the charging phase can be carried out through by driving the system to an excited state. The storage phase is detailed in this article. Finally, the QB discharging (i.e., the work extraction) may be performed by changing in time the couplings of the model, or, alternatively, by tuning the frequencies of the prescribed Hamiltonian.
@@ Acknowledgements.
We thank M. Polini and D. E. Chang for useful discussions and insightful suggestions. S.T. acknowledges financial support by MUR (Ministero dell’ Università e della Ricerca) through the following projects: PNRR MUR project PE0000023-NQSTI, PRIN 2017 Taming complexity via Quantum Strategies: a Hybrid Integrated Photonic approach (QUSHIP) Id. 2017SRN-BRK, and project PRO3 Quantum Pathfinder. G.M.A. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101002955 – CONQUER). G.C. acknowledges that results incorporated in this standard have received funding from the T-NiSQ consortium agreement financed by QUANTERA 2021 and that this work is supported in part by the Italian MUR Departments of Excellence grant 2023-2027 “Quantum Frontiers”.